OpenStudy (anonymous):

Question: Assume that the rate at which radioactive nuclei decay is proportional to the number of such nuclei that are present in a given sample. In a certain sample 10% of the orig. number of radioactive nuclei have undergone disintegration in a period of 100 yrs. a) what % of the orig. radioactive nuclei will remain after 1000 years I used the diff eq. dx/dt = -kx separating this to obtain x=x.e^(-kt) , where x. = orig. sample at the end I got 0.3486 or 34.86% and book says 59.05%, why????

OpenStudy (anonymous):

x=Ae^(-kt)

OpenStudy (anonymous):

0.9A=Ae^(-100k)

OpenStudy (anonymous):

from the first sentence.

OpenStudy (anonymous):

ln(0.9) = -100k k= 0.0010536....

OpenStudy (anonymous):

x = A e^(-kt) ----> restating the eqn

OpenStudy (anonymous):

when t=1000, x= A e^(-0.0010536 x 1000) = A x (0.34868...)

OpenStudy (anonymous):

You are correct, there is a mistake with the book.

OpenStudy (anonymous):

it still gives me 0.3486 and the book says 0.5905 :s I think the book is incorrect, then.